# DPMambaIR: All-in-One Image Restoration via Degradation-Aware Prompt State Space Model

### v1

@@ 3.3. DPMambaIR
We propose DPMambaIR, a degradation-aware framework for All-in-One image restoration. As illustrated in Figure 2, we first establish a Mamba-based image restoration baseline using an asymmetric encoder-decoder design within a U-shaped architecture. The restoration process begins with a $3\times 3$ convolutional layer, which extracts shallow features $\boldsymbol{F}\in\mathbb{R}^{H\times W\times C}$ from the degraded image. These features then pass through three encoding stages, each followed by downsampling. During each encoding stage, the resolution of the feature map is halved while the number of channels is doubled. Subsequently, the features pass through three decoding stages, where each stage involves upsampling and concatenating the output of the decoder with the corresponding output of the encoder to form the input for the next decoding stage. Finally, after a refinement stage, the clean image is restored through a shortcut connection that combines the input degraded image with the output of the decoder.
@@
3.4. Pretrained Degradation Extractor
Figure 3. (a) Degradation Modeling Process of Real-World Images, (b) The Architecture of Degradation Extractor.
Figure 4. The Architecture of (a) Degradation-aware Prompt Selective Scan (DPSS) and (b) High-frequency Enhancement Block (HEB), with Three Inputs $F_{1}$, $F_{2}$, and Degradation Embedding $D$.
To better extract degradation information from images, we revisit the degradation modeling process of real-world images, which spans from image capture and imaging to transmission, as illustrated in Figure  3 (a). This process typically involves four common types of degradations: (1) poor illumination conditions, (2) relative motion between the target objects and the camera, (3) artifacts introduced by transmission media in the optical path, and (4) compression during storage and transportation. We extend these four types of degradations to a more general framework and propose a unified degradation imaging model. The specific formulation is defined as follows:
(6) $I(x)=Q\big{(}(\alpha(x)J(x)+\beta(x)\gamma(x))\otimes K\big{)}.$
@@ 3.4. Pretrained Degradation Extractor
Based on the aforementioned degradation modeling process, we design a universal degradation extractor, as illustrated in Figure 3 (b). Specifically, we introduce central difference convolution(CDC)(Wang et al., 2023a; Peng et al., 2024a, b) to extract edge and texture degradation cues related to illumination and blur. In parallel, we employ three groups of CDC with different scales to capture artifacts of varying sizes, such as visual artifacts and snowflakes. Additionally, we integrate a traditional convolution branch to learn brightness-related representations.
@@
Furthermore, to train the degradation extractor with fine-grained degradation encoding capability, we propose a pretraining method for the degradation extractor based on degraded image reconstruction. This method leverages a hybrid dataset consisting of paired data with diverse types of degradation. The core of this process is training the degradation extractor $\mathcal{E}$ to extract fine-grained degradation information from the degraded image $\boldsymbol{I}_{D}$ and distill it into a compact degradation embedding $D$ of size $1\times 512$, formulated as $D=\mathcal{E}(\boldsymbol{I}_{D})$. Subsequently, the reconstruction unit $\mathcal{R}$ utilizes $D$ and the clean image $\boldsymbol{O}$ to reconstruct the degraded image $\boldsymbol{I}_{D}$, formulated as $\boldsymbol{\hat{I}}_{D}=\mathcal{R}(\boldsymbol{O},D)$.
@@
We further introduce a dual-component loss function $L$, which combines the $L_{1}$ norm and the Learned Perceptual Image Patch Similarity (LPIPS) metric. The combined loss function is expressed as $L(\boldsymbol{I}_{D},\boldsymbol{\hat{I}}_{D})=L_{1}+L_{LPIPS}$, striking a careful balance between reconstruction accuracy and perceptual quality.
3.5. Degradation-aware Prompt State Space Model
@@
To address these limitation mentioned in 3.2, we propose a novel Degradation-aware Prompt State Space Model to fully unlock the potential of SSM in All-in-One image restoration. Specifically, we propose a novel Degradation-aware Prompt Selective Scan, which incorporates degradation information into the State Space Model. This is achieved by leveraging a Degradation Embedding extracted from a pre-trained Degradation Extractor to provide supplementary degradation-related contextual information that the traditional SSM cannot capture. In addition, we modulate the parameter matrices in the state-space equations to establish a degradation-aware mechanism. This mechanism enables SSM to adapt to the imbalances caused by the inherent differences among sub-tasks in All-in-One image restoration. To further mitigate the issue of high-frequency information loss, which arises from balancing the differences among sub-tasks during training, we design a simple yet effective High-frequency Enhancement Block (HEB). This component ensures the preservation of high-frequency details and significantly enhances the overall performance of the model.
@@
Degradation-aware Prompt Selective Scan. We revisit the traditional SSM, whose parameter settings are shown in Equation (3). Matrices $A$, $B$, $C$, and $D$ control the SSM output. The state transition matrix $A$ compresses historical information, while the input matrix $B$ maps input signals to the state space, determining their influence on historical states. The output matrix $C$ maps historical states to the observable space, reflecting their impact on the output. Matrix $D$ provides a direct pathway from input to output, similar to a residual connection. In traditional SSM, these matrices and $\Delta$ are derived from:
(7) $A=init(),D=init(),$
@@
where $init()$ is an initialization method that does not depend on the input $x$. This way may lead to two key issues: (a) Traditional State Space Model struggles to effectively leverage global contextual information. (b) The lack of a proper degradation-aware mechanism makes it challenging to balance the differences among tasks in multi-task learning.
To address these issues, we first utilize a pre-trained degradation extractor $\mathcal{E}$ to extract the degradation embedding $D$ from the input image $I_{D}$. Subsequently, $D$ is used to modulate $\Delta$, the input matrix $B$, and the output matrix $C$ of the SSM. This introduces a degradation-aware mechanism in three aspects: historical information compression, input information processing, and global context awareness, as illustrated in Figure 4 (a). The specific method is formulated as follows:
(8) $D$ $=\mathcal{E}(I_{D}),$
$\alpha_{\Delta}$ $=\mathcal{M}_{\Delta}(D),\,\alpha_{B}=\mathcal{M}_{B}(D),\,\alpha_{C}=\mathcal{M}_{C}(D),$
@@
Here, $\mathcal{M}_{\Delta}$, $\mathcal{M}_{B}$, and $\mathcal{M}_{C}$ are combinations of linear layers used to map the degradation embedding $\mathbf{e}$ into feature spaces, forming modulation vectors $\alpha_{\Delta}$, $\alpha_{B}$, and $\alpha_{C}$ for $\Delta$, $B$, and $C$, respectively. $\Delta_{dp}$, $B_{dp}$, and $C_{dp}$ are the newly modulated SSM parameters. Finally, the state-space function formulation of the Degradation-aware State Space Model can be expressed as follows:
(9) $\overline{A}_{dp}$ $=\exp(\Delta_{dp}A),$
@@
Through this approach, the transformation matrix $A$ and input matrix $B$ gain degradation-aware capabilities, integrating degradation cues of various types and levels during historical information compression and input signal processing, making the model more task-friendly. The output matrix $C$ controls how historical information affects the output. In traditional SSM, the $i$-th pixel can only observe the $i-1$ preceding pixels, leading to a limited global receptive field. While existing methods address this with parallel multi-directional scanning, this increases computational costs and offers limited performance gains in low-level vision tasks (Guo et al., 2024c). Instead, we retain single-directional scanning and enhance global degradation cues via $C$, achieving efficient global awareness with lower computational overhead.
@@
Figure 5. We compare frequency-domain differences between clear and restored images under three configurations. The All-in-One setting shows a notable high-frequency gap compared to the Degradation-specific setting, which is significantly reduced by adding the High-frequency Enhancement Block (HEB), as shown by the blue line.
In All-in-One image restoration tasks, models tend to prioritize low-frequency restoration over high-frequency details, such as edges and textures, leading to their loss, as shown in Figure 5. To address this, we propose an efficient high-frequency enhancement block with minimal computational overhead. For input features $F$, we first extract channel-wise low-frequency features using global average pooling, then subtract them from $F$ to obtain high-frequency features, as shown in Figure 4(b). Finally, the high-frequency features are added back to enhance $F$, formulated as:
(10) ${F}_{enhanced}={F}+\alpha\cdot\big{(}{F}-\mathcal{G}({F})\big{)}.$
Here, $\alpha$ represents a learnable parameter that controls the weight of high-frequency information during the enhancement process. $\mathcal{G}$ represents the global average pooling operation. ${F}_{enhanced}$ denotes the features after high-frequency enhancement.
3.6. Loss Function
We employ the commonly used L1 loss $\mathcal{L}_{1}$ and L2 loss $\mathcal{L}_{2}$ in low-level vision tasks. Additionally, following previous works(Zamir et al., 2022; Cui et al., 2023), we incorporate a frequency-domain loss $\mathcal{L}_{fft}$ for training. The total loss is defined as:
(11) $\mathcal{L}_{total}=\lambda_{1}\cdot\mathcal{L}_{1}(O,\hat{O})+\lambda_{2}\cdot\mathcal{L}_{2}(O,\hat{O})+\lambda_{3}\cdot\mathcal{L}_{fft}(O,\hat{O}).$
@@ 3.6. Loss Function
4. Experiment And Analysis
@@ 4.1. Experimental Settings
Training Details. Building on previous works(Zamir et al., 2022; Potlapalli et al., 2023; Guo et al., 2024c), we develop a Mamba-based baseline for image restoration. Our method is trained using the AdamW(Kingma and Ba, 2014) optimizer with parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a weight decay of $1e^{-4}$ for 300,000 iterations. The initial learning rate of $3e^{-4}$ is gradually reduced to $1e^{-6}$ through cosine annealing. Following(Zamir et al., 2022), we employ a progressive training strategy, cropping image patches of size $192\times 192$. Horizontal and vertical flips are applied for data augmentation. Our experiments are conducted on a single NVIDIA RTX A100 GPU and implemented using the PyTorch platform.
@@
Datasets. For All-in-One image restoration, we collect a large-scale mixed dataset that includes seven classic and popular degradation types: fog, rain, snow, low light, blur, JPEG compression artifacts, and noise. Additionally, we evaluate performance on Degradation-specific tasks, using the LOL dataset(Wei et al., 2018) for low-light image enhancement, the GoPro dataset(Nah et al., 2017a) for image deblurring, the Rain100H dataset(Yang et al., 2017) for image deraining, and the RESIDE6K dataset(Qin et al., 2020) for image dehazing.
Table 1. Quantitative Comparison on Seven Datasets for All-in-One Image Restoration Task with State-of-the-Art Methods. The Best and Second-Best Results Are in Bold and Underline.
Method Lowlight Snow Rain Fog Jpeg Noise Blur Average PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM MPRNet 19.71 0.659 30.94 0.938 26.27 0.912 25.63 0.947 30.51 0.938 28.03 0.897 26.07 0.869 26.74 0.880 MIRNet 19.81 0.668 31.04 0.941 25.77 0.907 22.26 0.882 30.54 0.939 28.14 0.900 26.32 0.873 26.27 0.873 NAFNet 19.49 0.650 30.33 0.935 26.04 0.906 25.63 0.946 30.20 0.936 27.80 0.895 26.13 0.870 26.52 0.877 Restormer 19.99 0.670 31.67 0.944 26.91 0.919 26.18 0.953 30.56 0.939 28.17 0.890 26.62 0.880 27.16 0.886 PromptIR 19.90 0.669 31.52 0.943 26.80 0.918 26.13 0.945 30.57 0.939 28.18 0.901 26.36 0.874 27.07 0.884 IDR 19.68 0.663 31.16 0.941 26.39 0.914 26.11 0.948 30.51 0.939 28.10 0.899 26.29 0.873 26.89 0.883 OneRestore 19.77 0.657 29.95 0.931 25.37 0.900 27.74 0.962 30.10 0.934 27.65 0.890 25.90 0.866 26.64 0.877 AdaIR 19.99 0.672 31.94 0.923 27.01 0.922 26.61 0.952 30.53 0.939 28.19 0.900 26.61 0.879 27.26 0.887 DPMambaIR 20.04 0.680 32.36 0.949 27.07 0.926 28.13 0.963 30.57 0.939 28.25 0.902 27.43 0.894 27.69 0.893
Table 2. Quantitative Comparison on Deraining, Low-light Image Enhancement(LLIE), Deblurring and Dehazing for Degradation-Specific Image Restoration Task with State-of-the-Art Methods. The Best and Second-Best Results Are in Bold and Underline.
Deraining PSNR SSIM JORDER 26.25 0.835 PReNet 29.46 0.899 MPRNet 30.41 0.891 MAXIM 30.81 0.903 Restormer 31.46 0.904 DPMambaIR 32.30 0.967
LLIE PSNR SSIM EnlightenGAN 17.61 0.653 MIRNet 24.14 0.830 URetinex-Net 19.84 0.824 MAXIM 23.43 0.863 IAT 23.38 0.809 DPMambaIR 24.20 0.852
Deblurring PSNR SSIM DeepDeBlur 29.08 0.913 DeBlurGAN 28.70 0.858 DeBlurGANV2 29.55 0.934 MT-RNN 31.15 0.945 IR-SDE 30.70 0.901 DPMambaIR 31.42 0.948
Dehazing PSNR SSIM GCANet 26.59 0.935 GridDehazeNet 25.86 0.944 DeHazeFormer 30.29 0.964 MAXIM 29.12 0.932 DA-CLIP 30.16 0.936 DPMambaIR 30.89 0.977
@@
Figure 6. Qualitative comparison on Deblurring, Deraining and Dehazing for All-in-One Image Restoration Task.
@@
Evaluation Metrics. Following previous works(Zamir et al., 2022; Guo et al., 2024c), we use peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) to quantitatively compare different methods on the RGB channels.
@@
Comparisons with State-of-the-art Methods. For All-in-One image restoration, we compare our method against eight approaches: MIRNet (Zamir et al., 2020), NAFNet (Chen et al., 2022), MPRNet (Zamir et al., 2021), Restormer (Zamir et al., 2022), PromptIR (Potlapalli et al., 2023), IDR (Zhang et al., 2023a), OneRestore (Guo et al., 2024a), and AdaIR (Cui et al., 2025). Among these, MIRNet, NAFNet, MPRNet, and Restormer are general-purpose image restoration models, while the latter four methods are specifically designed for All-in-One image restoration. Specifically, PromptIR employs visual prompts to guide restoration, OneRestore leverages a pre-trained degradation category encoder with cross-attention mechanisms, and AdaIR utilizes frequency domain information for adaptive restoration. For Degradation-specific image restoration tasks, we compare our method with different state-of-the-art approaches tailored to each domain. The details are as follows: (1) For deraining, we compare against JORDER (Yang et al., 2019), PReNet (Ren et al., 2019), MPRNet (Zamir et al., 2021), MAXIM (Tu et al., 2022), and Restormer (Zamir et al., 2022). (2) For low-light image enhancement, we compare against EnlightenGAN (Jiang et al., 2021), MIRNet (Zamir et al., 2020), Uretinex-Net (Wu et al., 2022), MAXIM (Tu et al., 2022), and IAT (Cui et al., 2022). (3) For deblurring, we compare against DeepDeblur (Nah et al., 2017b), DeblurGAN (Kupyn et al., 2017), DeblurGANv2 (Kupyn et al., 2019), MT-RNN (Park et al., 2020), and IR-SDE (Luo et al., 2023b). (4) For dehazing, we compare against GCANet (Chen et al., 2018), GridDehazeNet (Liu et al., 2019), DeHazeFormer (Song et al., 2023), MAXIM (Tu et al., 2022), and DA-CLIP (Luo et al., 2023a).
4.2. Quantitative Results

### v2

@@ III-B DPMambaIR
We propose DPMambaIR, a degradation-aware framework for All-in-One image restoration. As illustrated in Fig. 2, we first establish a Mamba-based image restoration baseline using an asymmetric encoder-decoder design within a U-shaped architecture. The restoration process begins with a $3\times 3$ convolutional layer, which extracts shallow features $\boldsymbol{F}\in\mathbb{R}^{H\times W\times C}$ from the degraded image. These features then pass through three encoding stages, each followed by downsampling. During each encoding stage, the resolution of the feature map is halved while the number of channels is doubled. Subsequently, the features pass through three decoding stages, where each stage involves upsampling and concatenating the output of the decoder with the corresponding output of the encoder to form the input for the next decoding stage. Finally, after a refinement stage, the clean image is restored through a shortcut connection that combines the input degraded image with the output of the decoder.
@@
III-C Degradation Extractor
To better extract degradation information from images, we revisit the degradation modeling process of real-world images, which spans from image capture and imaging to transmission, as illustrated in Fig. 3 (a). This process typically involves four common types of degradations: (1) poor illumination conditions, (2) relative motion between the target objects and the camera, (3) artifacts introduced by transmission media in the optical path, and (4) compression during storage and transportation. We unify these scenarios into a generalized degradation formation model:
$I(x)=Q\big((\alpha(x)J(x)+\beta(x)\gamma(x))\otimes K\big).$ (5)
@@ III-C Degradation Extractor
Figure 4: The Architecture of (a) Degradation-aware Prompt Selective Scan (DPSS) and (b) High-frequency Enhancement Block (HEB), with Three Inputs $F_{1}$, $F_{2}$, and Degradation Embedding $\boldsymbol{E}_{d}$.
@@
Based on the aforementioned degradation modeling process, we design a universal degradation extractor, as illustrated in Fig. 3 (b). Specifically, we introduce central difference convolution (CDC)[36, 37, 38] to extract edge and texture degradation cues related to illumination and blur. In addition, we propose a multi-scale degradation perception module, which employs four parallel CDC operators of different sizes to capture degradation features at various scales. By concatenating the outputs, the module retains complete degradation representations across all frequency components. Furthermore, we employ three groups of CDC with different scales to capture artifacts of varying sizes, such as visual artifacts and snowflakes. Additionally, we integrate a traditional convolution branch to learn brightness-related representations.
To achieve a balance between extraction performance and computational cost, we introduce a re-parameterization technique for efficient integration of degradation representations, as illustrated in Fig. 3 (c). These two implementations correspond to different variants of our method, which are presented in Table I as DPMambaIR-L and DPMambaIR, respectively.
@@
Furthermore, to endow the degradation extractor with fine-grained degradation encoding capability, we propose a pretraining strategy based on degraded image reconstruction. This strategy leverages a hybrid dataset composed of paired samples with diverse degradation types. The key objective is to train the degradation extractor $\mathcal{E}$ to capture detailed degradation information from a degraded image $\boldsymbol{I}_{D}$ and compress it into a compact embedding $\boldsymbol{E}_{d}$, formulated as:
$\boldsymbol{E}_{d}=\mathcal{E}(\boldsymbol{I}_{D}).$ (6)
Subsequently, the reconstruction unit $\mathcal{R}$ takes $\boldsymbol{E}_{d}$ together with the clean content embedding $\boldsymbol{E}_{c}$ to reconstruct the degraded image $\boldsymbol{I}_{D}$. In particular, we adopt Restormer as the content extractor to derive $\boldsymbol{E}_{c}$ from the clean reference image $\boldsymbol{O}$, expressed as:
$\boldsymbol{E}_{c}=Restormer(\boldsymbol{O}),\ \ \boldsymbol{\hat{I}}_{D}=\mathcal{R}(\boldsymbol{E}_{c},\boldsymbol{E}_{d}).$ (7)
@@
This pretraining paradigm encourages the degradation extractor to encode both the type and severity of degradation into a continuous embedding space, which then serves as a compact and informative task descriptor for downstream restoration. We further introduce a dual-component loss function $L$, which combines the $L_{1}$ norm and the Learned Perceptual Image Patch Similarity (LPIPS) metric. The combined loss function is expressed as $L(\boldsymbol{I}_{D},\boldsymbol{\hat{I}}_{D})=L_{1}+L_{LPIPS}$, striking a balance between reconstruction accuracy and perceptual quality.
III-D Degradation-aware Prompt State Space Model
@@
To address the limitations of SSM-based image restoration—specifically the insufficient use of global context and the lack of effective degradation-aware mechanisms, we propose a novel Degradation-aware Prompt State Space Model (DP-SSM) to fully unlock the potential of SSM in All-in-One image restoration. Specifically, we propose a novel Degradation-aware Prompt Selective Scan, which incorporates degradation information into the State Space Model. This is achieved by leveraging a Degradation Embedding extracted from a pre-trained Degradation Extractor to provide supplementary degradation-related contextual information that the traditional SSM cannot capture. In addition, we modulate the parameter matrices in the state-space equations to establish a degradation-aware mechanism. This mechanism enables SSM to adapt to multi-task learning and handle task competition. To better recover local high-frequency details, we design a lightweight High-frequency Enhancement Block (HEB). This component ensures the preservation of high-frequency details and enhances the overall model performance.
@@
Degradation-aware Prompt Selective Scan. We revisit the traditional SSM, whose parameter settings are shown in Eq.(2). Matrices $A$, $B$, $C$, and $D$ control the SSM output. The state transition matrix $A$ compresses historical information, while the input matrix $B$ maps input signals to the state space, determining their influence on historical states. The output matrix $C$ maps historical states to the observable space, reflecting their impact on the output. Matrix $D$ provides a direct pathway from input to output, similar to a residual connection. In traditional SSM, these matrices and $\Delta$ are derived from:
$A=init(),D=init(),$ (8)
@@
where $init()$ is an initialization method that does not depend on the input $x$. However, such parameterization does not incorporate degradation information, which motivates us to introduce a degradation-aware prompt mechanism to guide the state-space modeling.
To integrate degradation awareness into the SSM, we first utilize a degradation extractor $\mathcal{E}$ to extract the degradation embedding $\boldsymbol{E}_{d}$ from the input image $I_{D}$. This embedding is then employed to modulate $\Delta$, the input matrix $B$, and the output matrix $C$, thereby enhancing the model’s ability to compress historical information, process input signals, and capture global context, as illustrated in Fig. 4 (a). The specific formulation is given as follows:
$\boldsymbol{E}_{d}$ $=\mathcal{E}(I_{D}),$ (9)
$\alpha_{\Delta}$ $=\mathcal{M}_{\Delta}(\boldsymbol{E}_{d}),\,\alpha_{B}=\mathcal{M}_{B}(\boldsymbol{E}_{d}),\,\alpha_{C}=\mathcal{M}_{C}(\boldsymbol{E}_{d}),$
@@
Here, $\mathcal{M}_{\Delta}$, $\mathcal{M}_{B}$, and $\mathcal{M}_{C}$ are combinations of linear layers used to map the degradation embedding $\boldsymbol{E}_{d}$ into feature spaces, forming modulation vectors $\alpha_{\Delta}$, $\alpha_{B}$, and $\alpha_{C}$ for $\Delta$, $B$, and $C$, respectively. $\Delta_{dp}$, $B_{dp}$, and $C_{dp}$ are the newly modulated SSM parameters.
Finally, the formulation of the Degradation-aware State Space Model can be expressed as follows:
$\overline{A}_{dp}$ $=\exp(\Delta_{dp}A),$ (10)
@@
Through this approach, the transformation matrix $A$ and input matrix $B$ gain degradation-aware capabilities, integrating degradation cues of various types and levels during historical information compression and input signal processing, making the model more task-friendly. The output matrix $C$ controls how historical information affects the output. In traditional SSM, the $i$-th pixel can only observe the i1i-1 preceding pixels, leading to a limited global receptive field. While existing methods address this with parallel multi-directional scanning, this increases computational costs and offers limited performance gains in low-level vision tasks [32]. Instead, we retain single-directional scanning and enhance global degradation cues via $C$, achieving efficient global awareness with lower computational overhead.
@@
Figure 5: We compare frequency-domain differences between clear and restored images under three configurations. The All-in-One setting shows a notable high-frequency gap compared to the Degradation-specific setting, which is significantly reduced by adding the High-frequency Enhancement Block (HEB), as shown by the blue line.
All-in-One image restoration involves recovering image components distributed across different frequency bands. While our proposed Degradation-aware Prompt State Space Model enables dynamic adaptation to degradation characteristics and partially alleviates inter-task conflicts, it still exhibits a bias toward low-frequency restoration due to the inherent global modeling nature of Mamba. As shown in Fig. 5, All-in-One models tend to underperform in reconstructing high-frequency details such as edges and textures, compared to degradation-specific models. To address this limitation, we introduce an extremely lightweight High-Frequency Enhancement Block (HEB) as a complementary module.
For input features $F$, we first extract channel-wise low-frequency features using global average pooling, then subtract them from $F$ to obtain high-frequency components, as illustrated in Fig. 4(b). Finally, the high-frequency features are added back to enhance $F$, formulated as:
${F}_{enhanced}={F}+\alpha\cdot\big({F}-\mathcal{G}({F})\big).$ (11)
Here, $\alpha$ represents a learnable scalar that controls the contribution of high-frequency enhancement, and $\mathcal{G}$ denotes the global average pooling operation. ${F}_{enhanced}$ is the enhanced feature map.
III-E Loss Function
We employ the commonly used L1 loss $\mathcal{L}_{1}$ and L2 loss $\mathcal{L}_{2}$ in low-level vision tasks. Additionally, following previous works [39, 40], we incorporate a frequency-domain loss $\mathcal{L}_{fft}$ for training. The total loss is defined as:
[…]
@@ III-E Loss Function
Figure 6: Examples of Seven Types of Image Degradations in the Dataset.
TABLE I: Quantitative Comparison on Seven Datasets for All-in-One Image Restoration Task with State-of-the-Art Methods. The Best and Second-Best Results Are in Bold and Underline.
Method Lowlight Snow Rain Haze Jpeg Noise Blur Average PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM MPRNet 19.71 0.659 30.94 0.938 26.27 0.912 25.63 0.947 30.51 0.938 28.03 0.897 26.07 0.869 26.74 0.880 MIRNet 19.81 0.668 31.04 0.941 25.77 0.907 22.26 0.882 30.54 0.939 28.14 0.900 26.32 0.873 26.27 0.873 NAFNet 19.49 0.650 30.33 0.935 26.04 0.906 25.63 0.946 30.20 0.936 27.80 0.895 26.13 0.870 26.52 0.877 Restormer 19.99 0.670 31.67 0.944 26.91 0.919 26.18 0.953 30.56 0.939 28.17 0.890 26.62 0.880 27.16 0.886 MambaIR 19.81 0.664 30.43 0.935 25.52 0.901 25.90 0.944 30.38 0.937 27.91 0.895 26.07 0.869 26.57 0.878 PromptIR 19.90 0.669 31.52 0.943 26.80 0.918 26.13 0.945 30.57 0.939 28.18 0.901 26.36 0.874 27.07 0.884 IDR 19.68 0.663 31.16 0.941 26.39 0.914 26.11 0.948 30.51 0.939 28.10 0.899 26.29 0.873 26.89 0.883 NDR-Restore 19.79 0.635 30.00 0.902 26.00 0.813 27.00 0.943 30.19 0.872 27.78 0.786 25.77 0.791 26.64 0.820 OneRestore 19.77 0.657 29.95 0.931 25.37 0.900 27.74 0.962 30.10 0.934 27.65 0.890 25.90 0.866 26.64 0.877 MoCEIR 19.92 0.674 31.22 0.943 26.56 0.920 28.06 0.963 30.51 0.939 28.13 0.900 26.34 0.874 27.25 0.888 AdaIR 19.99 0.672 31.94 0.923 27.01 0.922 26.61 0.952 30.53 0.939 28.19 0.900 26.61 0.879 27.26 0.887 DPMambaIR 20.04 0.680 32.36 0.949 27.07 0.926 28.13 0.963 30.57 0.939 28.25 0.902 27.43 0.894 27.69 0.893 DPMambaIR-L 20.06 0.677 32.31 0.950 27.11 0.925 28.39 0.963 30.55 0.939 28.21 0.902 27.45 0.895 27.73 0.893
TABLE II: Quantitative Comparison on Deraining, Low-light Image Enhancement(LLIE), Deblurring and Dehazing for Degradation-Specific Image Restoration Task with State-of-the-Art Methods. The Best and Second-Best Results Are in Bold and Underline.
Deraining PSNR SSIM JORDER 26.25 0.835 PReNet 29.46 0.899 MPRNet 30.41 0.891 MAXIM 30.81 0.903 Restormer 31.46 0.904 OneRestore 29.36 0.944 MoCEIR 31.23 0.960 AdaIR 31.46 0.961 DPMambaIR 32.30 0.967
LLIE PSNR SSIM EnlightenGAN 17.61 0.653 MIRNet 24.14 0.830 URetinex-Net 19.84 0.824 MAXIM 23.43 0.863 IAT 23.38 0.809 OneRestore 22.97 0.835 MoCEIR 23.64 0.818 AdaIR 23.47 0.831 DPMambaIR 24.20 0.852
Deblurring PSNR SSIM DeepDeBlur 29.08 0.913 DeBlurGAN 28.70 0.858 DeBlurGANV2 29.55 0.934 MT-RNN 31.15 0.945 IR-SDE 30.70 0.901 OneRestore 28.76 0.915 MoCEIR 30.05 0.933 AdaIR 30.57 0.939 DPMambaIR 31.42 0.948
Dehazing PSNR SSIM GCANet 26.59 0.935 GridDehazeNet 25.86 0.944 DeHazeFormer 30.29 0.964 MAXIM 29.12 0.932 DA-CLIP 30.16 0.936 OneRestore 30.16 0.974 MoCEIR 30.62 0.975 AdaIR 30.87 0.975 DPMambaIR 30.89 0.977
IV Experiment and Analysis
@@ IV-A Experimental Settings
Training Details. Building on previous works[39, 22, 32], we develop a Mamba-based baseline for image restoration. Our method is trained using the AdamW[41] optimizer with parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a weight decay of $1e^{-4}$ for 300,000 iterations. The initial learning rate of $3e^{-4}$ is gradually reduced to $1e^{-6}$ through cosine annealing. Following[39], we employ a progressive training strategy, cropping image patches of size $192\times 192$. Horizontal and vertical flips are applied for data augmentation. Our experiments are conducted on a single NVIDIA RTX A100 GPU and implemented using the PyTorch platform.
@@
Datasets. For All-in-One image restoration, we collect a large-scale mixed dataset covering seven common degradation types: haze, rain, snow, low-light, blur, noise, and JPEG compression, as shown in Fig. 6. Detailed information regarding these datasets is summarized in Table III. Additionally, degradation-specific evaluations are conducted using LOL [42] for low-light enhancement, GoPro [43] for deblurring, Rain100H [44] for deraining, and RESIDE6K [45] for dehazing.
TABLE III: Datasets for Image Restoration Tasks
Degradation Dataset Train Pairs Test Pairs
Haze RESIDE6K [45] 6000 1000
Rain Rain13K [46] 13711 -
Rain100H [44] - 100
Snow Snow100K [47] 5000 1000
Low Light LOL [42], LSRW [48] 6085 65
Blur GoPro [43] 2103 1111
Noise WaterlooED [49] 4744 -
BSD400 [50] - 400
JPEG WaterlooED [49] 4744 -
BSD400 [50] - 400
@@
Figure 7: Qualitative comparison on Deblurring, Deraining and Dehazing for All-in-One Image Restoration Task.
@@
Evaluation Metrics. Following previous works[39, 32], we use peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) to quantitatively compare different methods on the RGB channels.
@@
Comparisons with State-of-the-art Methods. For All-in-One image restoration, we compare against eleven methods: MIRNet [51], NAFNet [52], MPRNet [53], Restormer [39], MambaIR [32], PromptIR [22], IDR [23], NDR-Restore [54], OneRestore [9], MoCEIR [55], and AdaIR [24]. The first four are general-purpose restoration models, while the latter four are designed specifically for All-in-One restoration: PromptIR uses visual prompts, OneRestore employs a pre-trained degradation encoder with cross-attention, and AdaIR leverages frequency-domain information for adaptive restoration. For degradation-specific restoration, we also include representative single-task methods such as JORDER [56] (deraining), MAXIM [57] (deblurring), EnlightenGAN [58] (low-light enhancement), DeepDeblur [59], AdaIR [24], DA-CLIP [25], among others.
IV-B Quantitative Results
All-in-one Image Restoration. We conducted two versions of our method for comparison: the normal version (DPMambaIR) and a Large version (DPMambaIR-L). The normal version adopts the degradation extractor illustrated in Fig. 3(c), while the Large version uses the architecture shown in Fig. 3(b). For subsequent evaluations, we primarily adopt the normal version unless otherwise specified. We compare our method against eleven state-of-the-art image restoration approaches. As shown in Table I, for the All-in-One image restoration task, our method achieves the best PSNR and SSIM on the aforementioned mixed datasets, outperforming existing methods. Specifically, our method achieves a PSNR of 27.69dB and an SSIM of 0.893, surpassing AdaIR by 0.43dB and 0.011, respectively. Additionally, DPMambaIR-L achieves a PSNR of 27.73dB and an SSIM of 0.893, slightly outperforming the normal version.